a motor car is going due north at a speed of 50km per hour. it makes a

900 left turn without changing the speed. the change in the velocity of the car is about


1
Expert's answer
2019-12-16T11:05:30-0500

Mechanics | Relativity


We need to find the change in the velocity of the car .


Solution:


We can consider these velocities on an xy - plane with the vectors "\bar {i} \space and \space \bar {j}"


"\bar {i} = East"


"- \bar {i} = West"

"\bar {j} = north"

,


"- \bar {j} = South"

Initial Velocity = a motor car is going due north at a speed of 50km per hour.


"Initial \space Velocity = 50 \bar {j}"

Final velocity = it makes a 90 Degrees left turn without changing the speed

= It moves in the direction of west with the speed 50 km/h



"Final \space Velocity = - 50 \bar{i}"

Now,


Change in velocity (Δv) = Final velocity – Initial velocity = "- 50 \bar{i} - 50 \bar {j}"


"- \bar {i}" direction (west) and "- \bar {j}" direction (south), so the total change in velocity will be pointing towards the southwest.


"|\Delta V| = \sqrt {(-50)^2 + (-50)^2} = \sqrt {2500+2500} = \sqrt {5000} = 70.7 \space kmph"


Answer: Yes


Change in velocity = "- 50 \bar{i} - 50 \bar {j}"

"|\Delta V| = 70.7 \space kmph"